a person has 100 dollar to spend on two goods x and y whose respectively prices are 3 dollar and 5 dollar.

draw the budget line.



1
Expert's answer
2022-01-11T09:32:07-0500

Solution:

"p_x=3, G_x=\frac {100}{3}=33\frac{1}{3}""p_y=5, G_y=\frac{100}{5}=20"